# Rajamanickam, Almutari and Majumdar

### v2

@@ 3 Spatially homogeneous systems
Once $S$ is determined, $R$ is calculated from the sister relation
@@ 3.1 Limiting solutions for small and large values of $W$
Consider small non-zero values of $W$ so that we can compute an asymptotic expansion of the seven critical points of $f_{m}$, as a polynomial in $W$. Keeping $\pazocal{A}$ and other parameters fixed, with errors of order $W^{2}$,
@@
Turning to the limit $W\to\infty$ for fixed $\pazocal{A}$, the modified bulk potential in (10) simplifies, as a first approximation, to $CW[\tfrac{1}{2}\Sigma_{s}\mathrm{tr}\,\mathbf{Q}^{2}-\mathrm{tr}(\mathbf{Q}\mathbf{X})]$, which is the net homogenized NP-LC interaction energy density. This leading order energy is minimized uniquely by $\mathbf{Q}=\mathbf{X}/\Sigma_{s}$ or $(S,R)=(s_{*}/\Sigma_{s},r_{*}/\Sigma_{s})$. In this limiting case, the biaxiality parameter of the energy minimizer, $\mathbf{Q}$, is equal to that of $\mathbf{X}$, i.e., $\beta=\beta_{*}$. On the other hand, as $\pazocal{A}\to+\infty$ for a fixed $W$, $f_{m}$ is uniquely minimized by an almost isotropic phase, i.e., $\mathbf{Q}\to W\mathbf{X}/\pazocal{A}$. In other words, the qualitative features of the original bulk potential, $f_{B}$ in (5), are preserved in the $\pazocal{A}\to\pm\infty$ limits, for fixed $W$.
@@ 3.2 Critical points for isotropic $\mathbf{X}=0$
When the tensor $\mathbf{X}$ is isotropic, i.e., $\mathbf{X}=0$, the collective effect of the NPs is also isotropic and the effective temperature is translated from $\pazocal{A}$ to $\tilde{\pazocal{A}}=\pazocal{A}+\Sigma_{s}W$. In other words, the effective temperature increases in the presence of colloidal NPs when $W>0$. The critical points are then the same as the seven solutions mentioned in the beginning of Section 3.1 for $W=0$, except that the isotropic phase is stable for $\tilde{\pazocal{A}}>0$ or for $\pazocal{A}>-\Sigma_{s}W$; the isotropic-nematic transition occurs at $\tilde{\pazocal{A}}=\pazocal{B}^{2}/27$ or for $\pazocal{A}=\pazocal{B}^{2}/27-\Sigma_{s}W$ and the ordered LC states do not exist for $\pazocal{A}>\pazocal{B}^{2}/24-\Sigma_{s}W$. Thus, the colloidal NPs broaden the window of stability of the isotropic phase and shrink the domain of existence for the ordered LC states.
@@ 3.3.1 Prolate uniaxial $\mathbf{X}$
The three roots of this polynomial correspond to the first three uniaxial solutions, $S_{0}$ and $S_{\pm}^{x}$, listed in Table 1. The uniaxial solution, $(S_{0},0)$, which emerges from the isotropic branch, only exists when $\tilde{\pazocal{A}}<\tilde{\pazocal{A}}_{3}$ and is oblate when $\mathbf{X}$ is prolate, and vice versa. The discriminant of the cubic polynomial, $\Delta=9\tilde{\pazocal{A}}^{2}\pazocal{B}^{2}-12\tilde{s}_{*}\pazocal{B}^{3}-216\tilde{\pazocal{A}}^{3}+324\tilde{s}_{*}\tilde{\pazocal{A}}\pazocal{B}-972\tilde{s}_{*}^{2}$, shows that there are three roots for sufficiently large and negative values of $\tilde{\pazocal{A}}$. Two of the roots coalesce at a critical value $\tilde{\pazocal{A}}_{3}$, where $\Delta=0$, and there exists only one real root for $\tilde{\pazocal{A}}>\tilde{\pazocal{A}}_{3}$ or for $\pazocal{A}>\tilde{\pazocal{A}}_{3}-\Sigma_{s}W$. An illustrative plot of the three roots as a function of temperature, $\pazocal{A}=\tilde{\pazocal{A}}-\Sigma_{s}W$ is shown in Figure 1(a).
@@
In summary, there are seven distinct critical points of $f_{m}$, for low temperatures, including co-existing locally-stable biaxial and globally-stable uniaxial bulk phases. The locally stable biaxial phases are induced by the colloidal NPs and cannot be observed with $W=0$ or in the pure cholesteric system. The two biaxial local minimizers, $(S,R)=(S_{+}^{y},R_{+}^{y})$ and $(S,R)=(S_{+}^{z},R_{+}^{z})$, exist only for $\tilde{\pazocal{A}}<\tilde{\pazocal{A}}_{4}$, as demarcated in Figure 1. There is no isotropic-nematic/cholesteric phase transition as such, when $W\neq 0$ and $\mathbf{X}\neq 0$. At high temperatures, we only observe the globally stable uniaxial critical point, $(S,R)=(S_{+}^{x},0)$, where $S_{+}^{x}\to\tilde{s_{*}}/\pazocal{A}$ as $A\to+\infty$. Furthermore, the critical point, $(S,R)=(S_{+}^{x},0)$, is the global minimizer of the modified bulk potential for all temperatures.
@@
$(S_{+}^{x},0)$, uniaxial, prolate ($x$-axis) global minimum ($\tilde{\pazocal{A}}<\tilde{\pazocal{A}}_{3}$)
$(S_{-}^{x},0)$, uniaxial, oblate ($x$-axis) global minimum ($\tilde{\pazocal{A}}_{4}^{+}<\tilde{\pazocal{A}}$)
$(S_{+}^{y},R_{+}^{y})$, biaxial, prolate ($y$-axis) local minimum ($\tilde{\pazocal{A}}<\tilde{\pazocal{A}}_{3}$), global minimum ($\tilde{\pazocal{A}}_{3}<\tilde{\pazocal{A}}<\tilde{\pazocal{A}}_{4}^{+}$)
@@
$(S_{+}^{z},R_{+}^{z})$, biaxial, prolate ($z$-axis) local minimum ($\tilde{\pazocal{A}}<\tilde{\pazocal{A}}_{3}$), global minimum ($\tilde{\pazocal{A}}_{3}<\tilde{\pazocal{A}}<\tilde{\pazocal{A}}_{4}^{+}$)
@@ 3.3.2 Oblate uniaxial $\mathbf{X}$
When $s_{*}$ is negative in (25), the prescribed $\mathbf{X}$ tensor exhibits oblate symmetry about the $x$-axis. While the same cubic and quartic equations (26)-(28) still apply, the stability properties are significantly different from the prolate case. We still have three uniaxial critical points from the cubic polynomial: $(S_{0},0)$, $(S_{+}^{x},0)$, $(S_{-}^{x},0)$ such that $S_{+}^{x}$ is globally stable for $\pazocal{A}<\pazocal{A}_{3}$ and does not exist for $\pazocal{A}>\pazocal{A}_{3}$. This is in stark contrast to the case of prolate uniaxial $\mathbf{X}$, for which $(S_{+}^{x},0)$ is globally stable for all $\pazocal{A}$. The critical point, $(S,R)=(S_{0},0)$, which is the continuation of the isotropic branch, has prolate symmetry for oblate $\mathbf{X}$. The oblate uniaxial critical point, $(S_{-}^{x},0)$, is unstable for $\pazocal{A}<\pazocal{A}_{4}^{+}$, but becomes the unique globally stable critical point for high temperatures, i.e., for $\pazocal{A}>\pazocal{A}_{4}^{+}$. Therefore, there is an exchange of stability between prolate and oblate critical points at $\pazocal{A}=\pazocal{A}_{4}$. The locally stable prolate biaxial critical points, $(S_{+}^{y},R_{+}^{y})$ and $(S_{+}^{z},R_{+}^{z})$ exist only for $\pazocal{A}\leq\pazocal{A}_{4}$. The oblate biaxial critical points are always unstable. Notably, there are only stable biaxial phases in a narrow temperature range, $\pazocal{A}_{3}<\pazocal{A}<\pazocal{A}_{4}$, which is a novelty induced by the suspended NPs. As with the case of uniaxial prolate $\mathbf{X}$, the oblate uniaxial critical point approaches $S_{-}^{x}\to\tilde{s}_{*}/\pazocal{A}<0$ as $\pazocal{A}\to+\infty$ and there is no isotropic-nematic/cholesteric phase transition as such. However, there is co-existence between locally stable biaxial phases and the globally stable prolate uniaxial phase for $\pazocal{A}<\pazocal{A}_{3}$. The seven critical points are plotted in Figure 2 and their stability properties are summarised in Table 2.
@@ 4 Spatially inhomogeneous systems: Cholesteric-filled channel geometry
In Section 3, we focus on the bulk equilibrium phases determined by the minimizers and critical points of $f_{m}$ in (10). The homogenized bulk potential, $f_{m}$, does not distinguish between nematic and cholesteric liquid crystal phases [14]. The suspended NPs essentially break the symmetry of the nematic/cholesteric phase if $\mathbf{X}\neq 0$ and impose a preferred direction (or director) on homogeneous samples, for large $W$ or large $\pazocal{A}$.
@@
$\mathbf{Q}(0)=\mathbf{Q}(H)=\mathbf{Q}_{0}\equiv\begin{bmatrix}\tfrac{2}{3}s_{+}&0&0\\ 0&-\tfrac{1}{3}s_{+}&0\\ 0&0&-\tfrac{1}{3}s_{+}\end{bmatrix}$. (29)
@@
The equilibrium/physically observable cholesteric configurations are modelled by minimizers of the cholesteric free energy, subject to the imposed boundary conditions in (29). The effective cholesteric free energy is the sum of the elastic energy and the homogenized bulk energy, which captures the effects of the suspended NPs. The effective free energy per unit area is given by  [30, 13, 14]
@@ 5 Numerical results
Next, we present some heuristics for the case $\mathbf{X}\neq 0$. In the $\lambda\to\infty$ limit, the minimizers of (30) converge to the minimizer of the modified bulk potential $f_{m}$ in (10), almost everywhere [19]. The minimizers of $f_{m}$ (for this particular choice of $\mathbf{X}$) converge to the minimizers of $f_{B}$ as $\pazocal{A}\to$-$\infty$, for fixed $W$, i.e., uniaxial $\mathbf{Q}$ of the form (11) with arbitrary $\mathbf{n}$. In this case, informally speaking, we minimize the cholesteric free energy in the restricted class of $\mathbf{Q}$-tensors in (11) and recover the helical $\mathbf{Q}$-fields in (50). However, for modestly large values of $\pazocal{A}$ and fixed $W$, the minimizers instead converge to the uniaxial state described by
@@
in Figure 1, for sufficiently large values of $\lambda$. This motivates the untwisting transitions for large $\lambda$, as the temperature parameter $\pazocal{A}$ is increased, and we expect to see sharp transitions around $\pazocal{A}=-5$ since there is only one stable critical point, $(S_{+}^{x},0)$, of $f_{m}$ for $\pazocal{A}>-5$, as plotted in Figure 2.
@@
Figure 5: Numerical solutions of (33)-(37) subject to the Dirichlet conditions in (38) that demonstrate the transition from the uniaxial helical $\mathbf{Q}$-field to an uniaxial (with thin biaxial boundary layers at $z=0,1$) untwisted $\mathbf{Q}$-field, as $\pazocal{A}$ is increased. The parameter values are $\lambda=1000$ with $W=0.5$, $\pazocal{B}=1.5$, $\eta=4.5$, $\sigma=4\pi$, $\Sigma_{s}=4\pi$ (same parametric values are used in subsequent figures) and $\mathbf{X}=\left(\mathbf{e}_{x}\otimes\mathbf{e}_{x}-\frac{1}{3}\mathbf{I}\right)$. The two snippets on the left in each subfigure shows the dominant director field of $\mathbf{Q}$, projected onto xyxy- and yzyz-planes. The top row pertain to $\pi$-twisting helices and the bottom row to $2\pi$-twisting helices. In both cases, the helical $\mathbf{Q}$-fields disappear roughly around $\pazocal{A}=-5$ and this transition occurs at lower temperatures for larger $\lambda$.
@@
We numerically compute solutions of (33)-(37), subject to the Dirichlet boundary conditions (38), with $\mathbf{X}$ as in (49), and parametrically change $\pazocal{A}$ from $\pazocal{A}=-20$ to $\pazocal{A}=0$. The computations are carried out using COMSOL Multiphysics solvers. The results for $\lambda=1000$ are plotted in Figure 5 for the $\pi$-twisting and $2\pi$-twisting helices as initial guesses for the numerical solver. First of all, at a finite temperature (or at a finite $W$), the formula given in (50) is not strictly valid since a simple harmonic function is insufficient to characterise the solution, as can be inferred from the figure. However, the solutions are still found to be splay/bend free, except in small neighbourhoods of $z=0,1$. The plots for $\pazocal{A}=-6.6$ show that there is significant local biaxiality near the points $z_{n}=n/2\omega$, $n=1,3,5,\dots 2\omega-1$. Both the $\pi$-twisting and $2\pi$-twisting solutions then disappear roughly at $\pazocal{A}=-5$ and the solution transitions to the untwisted solution, dictated by the minimizer of $f_{m}$ for $\pazocal{A}=-5$ in Figure 1. Notably, this untwisting transition becomes more abrupt and occurs at lower temperatures for larger values of $\lambda$, where the bulk effects (governed by $f_{m}$) dominate the energy landscape. For instance, when $\lambda=7000$, the untwisting occurs near $\pazocal{A}=-8$. We also perform computations by decreasing $\pazocal{A}$ parametrically, using $\mathbf{Q}=\mathbf{X}/\Sigma_{s}$ as an initial guess. These suggest that the untwisted solution can persist to large negative values of $\pazocal{A}$, although we expect such states to be unstable at sufficiently low temperatures, due to the preference for twisted minimizers.
@@
Next, consider a much smaller domain corresponding to $\lambda=10$. We note that $\omega\pi$-twisting solutions with $\omega\neq 2$, are generally not energetically favourable for such small values of $\lambda$. In fact, the minimizer of the cholesteric free energy (30) converges to the biaxial, $2\pi$-twisting solution in  (43)-(47) in the $\lambda\to 0$ limit. The effect of changing $\pazocal{A}$ on these biaxial, $2\pi$-twisting solutions is illustrated in Figure 6. As $\pazocal{A}$ increases, the helical structure is gradually suppressed, although not completely eliminated in this regime. This indicates that the untwisting transition requires large positive values of $\pazocal{A}$ when $\lambda$ is small. In Figure 6, the asymptotic solutions, for the limits $\lambda\to 0$ with $\pazocal{A}\sim 1$ in Section 4.1, are plotted as dashed lines for the purpose of comparison. Although $\lambda=10$ is not a particularly small number, the agreement between the numerical solutions and the asymptotic solutions is satisfactory, especially for the value $\pazocal{A}=-1$.
@@
Figure 8: Numerical solutions of (33)-(37), subject to (38), with $\mathbf{X}=\left(\mathbf{e}_{z}\otimes\mathbf{e}_{z}-\frac{1}{3}\mathbf{I}\right)$ and $\lambda=1000$. These solutions illustrate the transition from the uniaxial, $2\pi$-twisting helical $\mathbf{Q}$-field to an uniaxial (with thin biaxial boundary layers at $z=0,1$) untwisted $\mathbf{Q}$-field, with increasing temperature.
@@
with $\pazocal{A}(0)=-20$ (or $\pazocal{A}(0)=0$), $\Delta\pazocal{A}=+0.5$ ($\Delta\pazocal{A}=-0.5$), $\Delta t=100$ and $\Delta t_{w}=10$. This function slowly increases (or decreases) $\pazocal{A}$ in steps of $\Delta A$ across every $\Delta t$ time interval; the jump in $\pazocal{A}$ is smoothed out by the $\tanh$ functions with a transition width of $\Delta t_{w}$. The quasi-steady variation allows us to assess the stability characteristics of the solutions in Figures 5-8, obtained by means of stationary computations. We find that the computed stationary solutions are stable. For large negative values of $\pazocal{A}$, the helical solutions are stable whereas the untwisted solutions are unstable. Moreover, when $\pazocal{A}$ is gradually increased from its initial value $\pazocal{A}(0)=-20$, the helical solutions transition to the untwisted solutions at values of $\pazocal{A}$ close to those found in stationary computations. Similarly, when $\pazocal{A}$ is decreased from $\pazocal{A}(0)=0$, the reverse transition (from untwisted to helical) occurs at lower temperatures than the helical-untwisted transition, thereby indicating the presence of hysteresis. Such hysteresis behaviour in cholesteric LC samples, doped with quantum dots, has been previously observed in the experiments conducted in [26]. This also opens avenues for changing the optical properties of a doped cholesteric system by means of increasing or decreasing temperature as opposed to applying external fields. In particular, we break the director symmetry with the $\mathbf{X}$-field–generated by a dilute suspension of colloidal NPs–rather than by an external electric field. This will be investigated further in future work.

### v3

@@ 3 Spatially homogeneous systems
[…]
@@ 3.1 Limiting solutions for small and large values of $W$
Consider small non-zero values of $W$ with $\mathbf{X}\neq 0$, so that we can compute an asymptotic expansion of the seven critical points of $f_{m}$, as a polynomial in $W$ (using the same notation for the critical points as for $W=0$). Keeping $\pazocal{A}$ and other parameters fixed, with errors of order $W^{2}$,
@@
Turning to the limit $W\to\infty$ for fixed $\pazocal{A}$, the modified bulk potential in (10) simplifies, as a first approximation, to $CW[\tfrac{1}{2}\Sigma_{s}\mathrm{tr}\,\mathbf{Q}^{2}-\mathrm{tr}(\mathbf{Q}\mathbf{X})]$, which is the net homogenized NP-LC interaction energy density. This leading order modified bulk energy is minimized uniquely by $\mathbf{Q}=\mathbf{X}/\Sigma_{s}$ or $(S,R)=(s_{*}/\Sigma_{s},r_{*}/\Sigma_{s})$. In this limiting case, the biaxiality parameter of the energy minimizer, $\mathbf{Q}$, is equal to that of $\mathbf{X}$, i.e., $\beta=\beta_{*}$. On the other hand, as $\pazocal{A}\to+\infty$ for a fixed $W$, $f_{m}$ is uniquely minimized by an almost isotropic phase, i.e., $\mathbf{Q}\to W\mathbf{X}/\pazocal{A}$. In other words, the qualitative features of the original bulk potential, $f_{B}$ in (5), are preserved in the $\pazocal{A}\to\pm\infty$ limits, for fixed $W$.
@@ 3.2 Critical points for isotropic $\mathbf{X}=0$
When the tensor $\mathbf{X}$ is isotropic, i.e., $\mathbf{X}=0$, the collective effect of the NPs is also isotropic and the effective temperature is translated from $\pazocal{A}$ to $\tilde{\pazocal{A}}=\pazocal{A}+\Sigma_{s}W$. In other words, the effective temperature increases in the presence of colloidal NPs when $W>0$. The critical points are then the same as the seven solutions mentioned at the beginning of Section 3.1 for $W=0$, except that the isotropic phase is stable for $\tilde{\pazocal{A}}>0$ or for $\pazocal{A}>-\Sigma_{s}W$; the isotropic-nematic transition occurs at $\tilde{\pazocal{A}}=\pazocal{B}^{2}/27$ or for $\pazocal{A}=\pazocal{B}^{2}/27-\Sigma_{s}W$ and the ordered LC states do not exist for $\pazocal{A}>\pazocal{B}^{2}/24-\Sigma_{s}W$. Thus, the colloidal NPs broaden the window of stability of the isotropic phase and shrink the domain of existence for the ordered LC states.
@@ 3.3.1 Prolate uniaxial $\mathbf{X}$
The three roots of this polynomial correspond to the first three uniaxial solutions, $S_{0}$ and $S_{\pm}^{x}$, listed in Table 1. The uniaxial solution, $(S_{0},0)$, which emerges from the isotropic branch, only exists when $\tilde{\pazocal{A}}<\tilde{\pazocal{A}}_{3}$ and is oblate when $\mathbf{X}$ is prolate, and vice versa. The discriminant of the cubic polynomial, $\Delta=9\tilde{\pazocal{A}}^{2}\pazocal{B}^{2}-12\tilde{s}_{*}\pazocal{B}^{3}-216\tilde{\pazocal{A}}^{3}+324\tilde{s}_{*}\tilde{\pazocal{A}}\pazocal{B}-972\tilde{s}_{*}^{2}$, shows that there are three roots for sufficiently large and negative values of $\tilde{\pazocal{A}}$ (when $\Delta>0$). Two of the roots coalesce at a critical value $\tilde{\pazocal{A}}_{3}$, where $\Delta=0$, and there exists only one real root for $\tilde{\pazocal{A}}>\tilde{\pazocal{A}}_{3}$ or for $\pazocal{A}>\tilde{\pazocal{A}}_{3}-\Sigma_{s}W$. An illustrative plot of the three roots as a function of temperature, $\pazocal{A}=\tilde{\pazocal{A}}-\Sigma_{s}W$ is shown in Figure 1(a).
@@
In summary, there are seven distinct critical points of $f_{m}$, for low temperatures, including co-existing locally stable biaxial and globally stable (global energy minimizers) uniaxial bulk phases. The locally stable biaxial phases are induced by the colloidal NPs and cannot be observed with $W=0$ or in the pure cholesteric system. The two biaxial local energy minimizers, $(S,R)=(S_{+}^{y},R_{+}^{y})$ and $(S,R)=(S_{+}^{z},R_{+}^{z})$, exist only for $\tilde{\pazocal{A}}<\tilde{\pazocal{A}}_{4}^{+}$, as demarcated in Figure 1. There is no isotropic-nematic/cholesteric phase transition as such, when $W\neq 0$ and $\mathbf{X}\neq 0$. At high temperatures, we only observe the globally stable uniaxial critical point, $(S,R)=(S_{+}^{x},0)$, where $S_{+}^{x}\to\tilde{s_{*}}/\pazocal{A}$ as $A\to+\infty$. Furthermore, the critical point, $(S,R)=(S_{+}^{x},0)$, is the global minimizer of the modified bulk potential for all temperatures.
@@
$(S_{+}^{x},0)$, uniaxial, prolate ($x$-axis) local minimum ($\tilde{\pazocal{A}}<\tilde{\pazocal{A}}_{3}$)
$(S_{-}^{x},0)$, uniaxial, oblate ($x$-axis) global minimum ($\tilde{\pazocal{A}}>\tilde{\pazocal{A}}_{4}^{+}$)
$(S_{+}^{y},R_{+}^{y})$, biaxial, prolate ($y$-axis) global minimum ($\tilde{\pazocal{A}}<\tilde{\pazocal{A}}_{4}^{+}$)
@@
$(S_{+}^{z},R_{+}^{z})$, biaxial, prolate ($z$-axis) global minimum ($\tilde{\pazocal{A}}<\tilde{\pazocal{A}}_{4}^{+}$)
@@ 3.3.2 Oblate uniaxial $\mathbf{X}$
When $s_{*}$ is negative in (25), the prescribed $\mathbf{X}$ tensor exhibits oblate symmetry about the $x$-axis. While the same cubic and quartic equations (26)-(28) still apply, the stability properties are significantly different from the prolate case. We still have three uniaxial critical points from the cubic polynomial: $(S_{0},0)$, $(S_{+}^{x},0)$, $(S_{-}^{x},0)$ such that $S_{+}^{x}$ is locally stable for $\pazocal{A}<\pazocal{A}_{3}$ and does not exist for $\pazocal{A}>\pazocal{A}_{3}$. This is in stark contrast to the case of prolate uniaxial $\mathbf{X}$, for which $(S_{+}^{x},0)$ is globally stable for all $\pazocal{A}$. The critical point, $(S,R)=(S_{0},0)$, which is the continuation of the isotropic branch, has prolate symmetry for oblate $\mathbf{X}$. The oblate uniaxial critical point, $(S_{-}^{x},0)$, is unstable for $\pazocal{A}<\pazocal{A}_{4}^{+}$ but becomes the unique globally stable critical point for high temperatures, i.e., for $\pazocal{A}>\pazocal{A}_{4}^{+}$. Therefore, there is an exchange of stability between prolate and oblate critical points at $\pazocal{A}=\pazocal{A}_{4}^{+}$. The globally stable prolate biaxial critical points, $(S_{+}^{y},R_{+}^{y})$ and $(S_{+}^{z},R_{+}^{z})$ exist only for $\pazocal{A}\leq\pazocal{A}_{4}^{+}$. The oblate biaxial critical points are always unstable. Notably, there are only stable biaxial phases in a narrow temperature range, $\pazocal{A}_{3}<\pazocal{A}<\pazocal{A}_{4}^{+}$, which is a novelty induced by the suspended NPs. As with the case of uniaxial prolate $\mathbf{X}$, the oblate uniaxial critical point approaches $S_{-}^{x}\to\tilde{s}_{*}/\pazocal{A}<0$ as $\pazocal{A}\to+\infty$ and there is no isotropic-nematic/cholesteric phase transition as such. However, there is co-existence between globally stable biaxial phases and the locally stable prolate uniaxial phase for $\pazocal{A}<\pazocal{A}_{3}$. The seven critical points are plotted in Figure 2 and their stability properties are summarised in Table 2.
@@ 4 Spatially inhomogeneous systems: Cholesteric-filled channel geometry
In Section 3, we focus on the bulk equilibrium phases determined by the minimizers and critical points of $f_{m}$ in (10). The modified bulk potential, $f_{m}$, does not distinguish between nematic and cholesteric liquid crystal phases [13]. The suspended NPs essentially break the symmetry of the nematic/cholesteric phase if $\mathbf{X}\neq 0$ and impose a preferred direction (or director) on homogeneous samples, for large $W$ or large $\pazocal{A}$.
@@
$\mathbf{Q}(0)=\mathbf{Q}(H)=\mathbf{Q}_{0}\equiv\begin{bmatrix}\tfrac{2}{3}s_{+}&0&0\\ 0&-\tfrac{1}{3}s_{+}&0\\ 0&0&-\tfrac{1}{3}s_{+}\end{bmatrix}$ (29)
@@
The equilibrium/physically observable cholesteric configurations are modelled by minimizers of the cholesteric free energy, subject to the imposed boundary conditions in (29). We take the effective cholesteric free energy to be the sum of the elastic energy and the modified bulk energy, which captures the effects of the suspended NPs. This is a formal approximation wherein we formally replace the collective effects of the suspended NPs by the modified bulk energy without any rigorous convergence analysis, in the dilute limit. The effective free energy per unit area is given by  [30, 12, 13]
@@ 5 Numerical results
Next, we present some heuristics for the case $\mathbf{X}\neq 0$. In the $\lambda\to\infty$ limit, the minimizers of (4) converge to the minimizer of the modified bulk potential $f_{m}$ in (10), almost everywhere [18]. The minimizers of $f_{m}$ converge to the minimizers of $f_{B}$ as $\pazocal{A}\to$-$\infty$, for fixed $W$, i.e., uniaxial $\mathbf{Q}$ of the form (11) with arbitrary $\mathbf{n}$. In this case, informally speaking, we minimize the cholesteric free energy in the restricted class of $\mathbf{Q}$-tensors in (11) and recover the helical $\mathbf{Q}$-fields in (50). However, for modestly large values of $\pazocal{A}$ and fixed $W$, the minimizers of $f_{m}$ (for the particular choice of $\mathbf{X}$ in (49); also see Figure 1) instead converge to the uniaxial state described by
@@
for sufficiently large values of $\lambda$. This motivates the untwisting transitions for large $\lambda$, as the temperature parameter $\pazocal{A}$ is increased, and we expect to see sharp transitions around $\pazocal{A}=-5$ since there is only one stable critical point, $(S_{+}^{x},0)$, of $f_{m}$ for $\pazocal{A}>-5$, as plotted in Figure 1.
@@
Figure 5: Numerical solutions of (33)-(37) subject to the Dirichlet conditions in (38) that demonstrate the transition from the uniaxial helical $\mathbf{Q}$-field to an uniaxial (with thin biaxial boundary layers at $z=0,1$) untwisted $\mathbf{Q}$-field, as $\pazocal{A}$ is increased. The parameter values are $\lambda=1000$ with $W=0.5$, $\pazocal{B}=1.5$, $\eta=4.5$, $\sigma=4\pi$, $\Sigma_{s}=4\pi$ (same parametric values are used in subsequent figures) and $\mathbf{X}=\left(\mathbf{e}_{x}\otimes\mathbf{e}_{x}-\frac{1}{3}\mathbf{I}\right)$. The two snippets on the left in each subfigure shows the dominant director field of $\mathbf{Q}$, projected onto xzxz- and yzyz-planes. The top row pertain to $\pi$-twisting helices and the bottom row to $2\pi$-twisting helices. In both cases, the helical $\mathbf{Q}$-fields disappear roughly around $\pazocal{A}=-5$ and this transition occurs at lower temperatures for larger $\lambda$.
@@
We numerically compute solutions of (33)-(37), subject to the Dirichlet boundary conditions (38), with $\mathbf{X}$ as in (49), and parametrically change $\pazocal{A}$ from $\pazocal{A}=-20$ to $\pazocal{A}=0$. The computations are carried out using COMSOL Multiphysics, which employs the finite element method for spatial discretization and the MUMPS (Multifrontal Massively Parallel Sparse) direct solver. A mesh size of $\Delta z=10^{-3}$ is used after verifying mesh independence. The results for $\lambda=1000$ are plotted in Figure 5 for the $\pi$-twisting and $2\pi$-twisting helices as initial guesses for the numerical solver. First of all, at a finite temperature (or at a finite $W$), the formula given in (50) is not strictly valid since a simple harmonic function is insufficient to characterise the solution, as can be inferred from the figure. However, the solutions are still found to be splay/bend free, except in small neighbourhoods of $z=0,1$. The plots for $\pazocal{A}=-6.6$ show that there is significant local biaxiality near the points $z_{n}=n/2\omega$, $n=1,3,5,\dots 2\omega-1$. Both the $\pi$-twisting and $2\pi$-twisting solutions then disappear roughly at $\pazocal{A}=-5$ and the solution transitions to the untwisted solution, dictated by the minimizer of $f_{m}$ for $\pazocal{A}=-5$ in Figure 1. Notably, this untwisting transition becomes more abrupt and occurs at lower temperatures for larger values of $\lambda$, where the bulk effects (governed by $f_{m}$) dominate the energy landscape. For instance, when $\lambda=7000$, the untwisting occurs near $\pazocal{A}=-8$. We also perform computations by decreasing $\pazocal{A}$ parametrically, using $\mathbf{Q}=\mathbf{X}/\Sigma_{s}$ as an initial guess. These suggest that the untwisted solution can persist to large negative values of $\pazocal{A}$, although we expect such solutions to be unstable at sufficiently low temperatures, due to the preference for twisted minimizers.
@@
Next, consider a much smaller domain corresponding to $\lambda=10$. We note that $\omega\pi$-twisting solutions with $\omega\neq 2$, are generally not energetically favourable for such small values of $\lambda$. In fact, the minimizer of the cholesteric free energy (4) converges to the biaxial, $2\pi$-twisting solution in  (43)-(47) in the $\lambda\to 0$ limit. The effect of changing $\pazocal{A}$ on these biaxial, $2\pi$-twisting solutions is illustrated in Figure 6. As $\pazocal{A}$ increases, the helical structure is gradually suppressed, although not completely eliminated in this regime. This indicates that the untwisting transition occurs for large positive values of $\pazocal{A}$, when $\lambda$ is small. In Figure 6, the asymptotic solutions for the limits $\lambda\to 0$ with $\pazocal{A}\sim 1$ in Section 4.1, are plotted as dashed lines for the purpose of comparison. Although $\lambda=10$ is not a particularly small number, the agreement between the numerical solutions and the asymptotic solutions is satisfactory, especially for the value $\pazocal{A}=-1$.
@@
Figure 8: Numerical solutions of (33)-(37), subject to (38), with $\mathbf{X}=\left(\mathbf{e}_{z}\otimes\mathbf{e}_{z}-\frac{1}{3}\mathbf{I}\right)$ and $\lambda=1000$. These solutions illustrate the transition from the uniaxial, $2\pi$-twisting helical $\mathbf{Q}$-field to an uniaxial (with thin biaxial boundary layers at $z=0,1$) untwisted $\mathbf{Q}$-field, with increasing temperature. In the last case, the leading eigenvector is $\mathbf{e}_{z}$ almost everywhere and therefore its projections onto the $xz$- and yzyz-planes are null.
@@
with $\pazocal{A}(0)=-20$ (or $\pazocal{A}(0)=0$), $\Delta\pazocal{A}=+0.5$ ($\Delta\pazocal{A}=-0.5$), $\Delta t=100$ and $\Delta t_{w}=10$. For time-dependent simulations, a relative tolerance of $10^{-3}$ and a fixed time step of $0.05$ are used in COMSOL Multiphysics. This function slowly increases (or decreases) $\pazocal{A}$ in steps of $\Delta A$ across every $\Delta t$ time interval; the jump in $\pazocal{A}$ is smoothed out by the $\tanh$ functions with a transition width of $\Delta t_{w}$. The quasi-steady variation allows us to assess the stability characteristics of the solutions in Figures 5-8, obtained by means of stationary computations. We find that the computed stationary solutions are stable. For large negative values of $\pazocal{A}$, the helical solutions are stable whereas the untwisted solutions are unstable. Moreover, when $\pazocal{A}$ is gradually increased from its initial value $\pazocal{A}(0)=-20$, the helical solutions transition to the untwisted solutions at values of $\pazocal{A}$ close to those found in stationary computations. Similarly, when $\pazocal{A}$ is decreased from $\pazocal{A}(0)=0$, the reverse transition (from untwisted to helical) occurs at lower temperatures than the helical-untwisted transition, thereby indicating the presence of hysteresis. Such hysteresis behaviour in cholesteric LC samples, doped with quantum dots, has been previously observed in the experiments conducted in [25]. This also opens avenues for changing the optical properties of a doped cholesteric system by means of increasing or decreasing temperature as opposed to applying external fields. In particular, we break the director symmetry with the $\mathbf{X}$-field–generated by a dilute suspension of colloidal NPs–rather than by an external electric field. This will be investigated further in future work.